a certain gas is compressed reversibly from 1 bar , 17 degrees to a pressure of 5bar in a perfectly thermally insulated cylinder the final temperature being 77degrees . The work done on the gas during the compression is 45kj/Kg. Calculate Cv alpha and R


1
Expert's answer
2021-05-20T20:13:27-0400

For reversible process"(\frac{P_1}{P_2})^n=\frac{T_1}{T_2}"

"n ln(\frac{P_1}{P_2})=\frac{T_1}{T_2}"

"n=(\gamma-1)/(\gamma)"

"W=R(T_1-T_2)/(\gamma-1)"

Put the Value of data given

"n ln(1\/5)=(290\/350)"

"n=0.1125"

Again,

"n=(\gamma-1)/\gamma"

"0.1125=(\gamma-1)/(\gamma)"

"\gamma=1.132"

And by work formula

"W=R(T_1-T_2)/(\gamma-1)"

"-45=R(290-350)\/(1.1132-1)"

R=0.099KJ/Kg-K

Also "C_v=\frac{R}{(\gamma-1)}"

"C_v=\frac{R}{(1.132-1)}"

Cv=0.75KJ/Kg-K